0 ts 20

a) State without any calculations which object has a greater displacement over the time interval shown in the graphs? Explain your answer

b) Calculate the displacement values for the objects A and B at the time moments of

- 10s

- 30s

- 60s

c) Assuming that at t=0 both objects were located at x= 0, make graphs of the position versus time for each object, for the entire interval from 0-60s. ( Don't have to actually do this but please explain how I do this)

d) Explain how yoy were able to make the graphs for the entire time interval having just 3 values of the displacement for each object
Answer #1


parta)

A)

D = (4*30)+(1*30) = 120+30 = 150 m

B)


D = (2*30)+(-2*30) = 0


DA > Db